. Rodney Cashman has been investing $2,000 quarterly for the past 18 years in an equity mutual fund. How much is the fund worth now assuming he has earned 11.5% compounded quarterly on his investment?    

N

I/Y

PV

PMT

FV
Answer #1

N =18*4 =72
I/Y =11.5%/4 =2.875%
PMT =2000
FV using formula
FV =PMT*((1+r)^n-1)/r =2000*((1+2.875%)^72-1)/2.875%)=465862.95

FV using financial calculator
N=72;I/Y =2.875%;PMT =-2000; CPT FV
FV =465862.95